### v1

# Transientangelo: Few-Viewpoint Surface Reconstruction Using Single-Photon Lidar

Abstract
We consider the problem of few-viewpoint 3D surface reconstruction using raw measurements from a lidar system. Lidar captures 3D scene geometry by emitting pulses of light to a target and recording the speed-of-light time delay of the reflected light. However, conventional lidar systems do not output the raw, captured waveforms of backscattered light; instead, they pre-process these data into a 3D point cloud. Since this procedure typically does not accurately model the noise statistics of the system, exploit spatial priors, or incorporate information about downstream tasks, it ultimately discards useful information that is encoded in raw measurements of backscattered light. Here, we propose to leverage raw measurements captured with a single-photon lidar system from multiple viewpoints to optimize a neural surface representation of a scene. The measurements consist of time-resolved photon count histograms, or transients, which capture information about backscattered light at picosecond time scales. Additionally, we develop new regularization strategies that improve robustness to photon noise, enabling accurate surface reconstruction with as few as 10 photons per pixel. Our method outperforms other techniques for few-viewpoint 3D reconstruction based on depth maps, point clouds, or conventional lidar as demonstrated in simulation and with captured data.
Figure 1: Transientangelo takes as input raw lidar scans from sparse viewpoints. These scans are used to optimize an SDF-based scene representation, which is further regularized to constrain the geometry from both captured and unseen viewpoints. The method recovers higher-fidelity surfaces than previous methods in the sparse-view and low-photon regime (i.e., from tens to hundreds of measured photons per pixel). The above scene was trained using simulated single-photon lidar data from five viewpoints, with an average of 150 photons per pixel over the occupied regions of the scene.
1 Introduction
Within the last few years, significant progress has been made on the problem of multi-view 3D surface reconstruction. Recent methods combine efficient neural surface representations with volumetric rendering models, enabling high-fidelity scene reconstruction from multi-view images [24, 48]. Overall, achieving accurate scene reconstruction is now relatively straightforward when tens to hundreds of input images are available. However, a significant remaining challenge is achieving robust reconstruction when the number of input viewpoints is reduced to only a few (e.g., less than ten). In this case, scene reconstruction from conventional images alone is ill-posed [19] and requires strong priors to solve robustly [55].
Instead of using conventional images, we investigate this challenge using lidar measurements captured from a few (2–5) different viewpoints. Lidar measures the distance to each scene point by illuminating it with a pulse of light and measuring the speed-of-light time delay of the back-reflected pulse. These measurements provide additional constraints to the multi-view reconstruction problem. However, conventional lidar systems do not output the raw, captured waveforms of backscattered light; instead, they post-process these data into a 3D point cloud. Since this procedure typically does not accurately model the noise statistics of the system, exploit spatial priors, or incorporate information about downstream tasks, it ultimately discards useful information that is encoded in raw measurements of backscattered light. We are interested in using the raw measurements captured by a single-photon lidar system—the time-resolved photon counts histograms, or transients, which contain information about the precise arrival times of backscattered light particles at picosecond time scales.
Previous methods for multi-viewpoint 3D reconstruction—whether based on conventional images or lidar—have several drawbacks. First, image-based methods are usually trained with tens or hundreds of images, making acquisition difficult [35, 32, 51, 24]. While several recent works attempt to tackle the few-view reconstruction problem [36, 29, 50, 63], they typically fail for the case where the baseline between input images becomes large and the problem becomes increasingly ill-posed. Recent methods that leverage large-scale diffusion models [55, 10] alleviate this limitation to some extent, but are also prone to hallucinating scene geometry. Other techniques improve few-view 3D reconstruction by introducing additional geometric constraints from point clouds [45] or monocular depth estimators [63]. In this vein, our work is closest to TransientNeRF [29], which shows that using raw lidar data instead of depth maps or point clouds leads to significant improvements in reconstructed scene appearance in the few-view setting. However, TransientNeRF is based on a volumetric representation and is designed for appearance modeling rather than surface reconstruction. Further, TransientNeRF uses long acquisition times ($\approx$20 minutes per view) to capture per-pixel photon count histograms with several thousand photon arrivals each.
To address these limitations we propose Transientangelo, a method for few-viewpoint scene reconstruction using transient supervision and a scene representation based on signed distance functions (SDFs). Inspired by Neuralangelo [24], we use an efficient neural representation with a hashing-based feature grid to parameterize an SDF and radiance field. The representation is then trained using supervision from photon count histograms. To further constrain the reconstructed surfaces, we regularize the scene geometry from unseen viewpoints and use supervision from the captured reflectivity (i.e., the time-integrated lidar measurements).
Overall, we improve few-viewpoint 3D surface reconstruction from raw lidar measurements compared to previous work (see Figure 1), while also demonstrating significant improvements in reconstructed geometry compared to image-based or depth-based approaches. Our approach also improves robustness to lidar measurements captured with relatively few photon counts (e.g., 10–300), which is especially important when operating at fast acquisition speeds or when imaging targets at long range. We make the following technical contributions:
  • We introduce Transientangelo: a method to perform surface reconstruction using few-viewpoint single-photon lidar.
  • We propose regularization techniques, that improve the reconstructed geometry and provide improved robustness in the low-flux regime where only tens to hundreds of photons are collected per pixel.
  • We introduce a multi-viewpoint simulated and captured transient dataset with a range of photon count levels, which we use to demonstrate state-of-the-art performance for few-viewpoint surface reconstruction.
2 Related Works
Figure 2: Transientangelo training procedure. For a pixel $\mathbf{p}$, we cast out a ray $\mathbf{r}(t)=\mathbf{o}+tc\bm{\omega}(\mathbf{p})$. Captured viewpoint: 3D ray coordinates are passed through the neural surface representation $\mathcal{F}$ to retrieve a radiance and an SDF value, which gets converted to density $\sigma$ (see Equation  3). These values are then binned into a transient, which after a convolution with the laser pulse gives the final rendered transient $\bm{\tau}_{f}[i,j,k]$. The network is supervised with an L1 loss between the rendered and captured transient and an L1 loss between the integrals of the rendered and captured transients (reflectivity loss). Unseen viewpoint: 3D ray coordinates are passed through the neural representation to retrieve rendering weights, which are used to calculate the variance of the weights around the depth $d$. The network is trained to minimize this variance, resulting in thinner surfaces and removing spurious zero-level sets.
Transientangelo brings together the areas of active single-photon imaging, neural surface reconstruction, and few-viewpoint 3D reconstruction.
Active single-photon imaging.
We are specifically concerned with lidar systems based on single-photon avalanche diodes (SPADs), which output precise timestamps corresponding to the arrival times of individual captured photons. For SPAD-based lidar, a transient is measured by repeatedly illuminating a point with pulses of light and accumulating the individual photon arrival times into a time-resolved histogram. Such captured histograms of a scene can be used for direct scene reconstruction [25, 15, 40, 29], non-line of sight scene reconstruction [3, 39, 44, 56, 7], seeing through participating media [26], visualizations of light propagation [28, 49, 11] or to uncover statistical properties of captured photons [42, 43]. Our work differs from these in that we specifically aim to recover high-fidelity surfaces from multi-viewpoint SPAD data. We are motivated by previous work[29] that demonstrates the strong geometric information encoded in transients and their promise for multi-viewpoint surface reconstruction.
While many previous techniques on 3D imaging with SPADs use high-powered lasers or long acquisition times to capture hundreds or thousands of photons [29, 25, 26, 39], our work is more closely related to single-photon imaging techniques that operate in the photon-starved regime. Such methods exploit spatial or temporal correlations in the photon arrivals to recover depth or reflectivity from only a few captured photons at each pixel [21, 42]. Our approach exploits cross-viewpoint information and reflectivity-based regularization for scene reconstruction from as few as 10 collected photons per pixel.
Neural surface reconstruction.
Neural surface representations typically use either occupancy [38, 37, 40] or a signed distance function (SDF) [51, 52, 24, 60, 61] to model scene geometry. The former uses a neural network to classify points in 3D space that fall inside an object [38]. The latter learns a network that predicts the SDF of a 3D point (i.e., its signed distance to the closest surface). The surface is extracted from the SDF by finding the zero-level set. Our work, inspired by Neuralangelo [24], builds on a hashgrid-based SDF representation  [35]. However, different from previous work, we use transient measurements as supervision.
Recent works [58, 14, 41, 9, 33] have also explored training neural surface representations using data from emerging sensors. Concurrent work [34] reconstructed 3D shapes of objects using transient measurements from a low-cost SPAD sensor. Their setup differs from ours as they utilize hundreds of training poses, while we investigate 3D surface reconstruction from 2 to 5 viewpoints. To the best of our knowledge, our work is the first to tackle the problem of few-viewpoint surface reconstruction using transient measurements.
Few-viewpoint neural reconstruction.
Prior work has addressed few-viewpoint reconstruction using geometric supervision (e.g., based on depth maps, point clouds, or transients) [29, 50, 63, 4, 46], or by introducing other strategies to regularize reconstruction from multi-view images [36, 63, 17, 20, 57, 62, 54, 59]. Our work is closest to TransientNeRF [29] and RegNeRF [36]. In contrast to TransientNeRF—which reconstructs a volumetric scene model from few-viewpoint transients—we focus on surface reconstruction using an SDF-based representation and improve robustness to noisy (low photon) measurements. Similar to RegNeRF, we regularize views rendered from novel viewpoints; however, we use a different depth-based regularization strategy and achieve improved performance in the few-view setting by exploiting transient data.
3 Method
In this section, we provide a mathematical description of single-photon lidar [29], which we then relate to the rendering model used to optimize the surface representation. An overview of the method is shown in Figure 2.
3.1 Measurement Model
Suppose that light from a laser pulse is reflected from a scene and focused onto a sensor at point $\mathbf{p}\in\mathbb{R}^{2}$. Also, assume that laser and sensor are co-axial, such that light propagates to the scene and back along the same path. Then, we can model the forward path of light along a ray $\mathbf{r}(t)=\mathbf{o}+tc\bm{\omega}(\mathbf{p})$, where $\mathbf{o}\in\mathbb{R}^{3}$ represents the camera origin, $c$ is the speed of light, and $\bm{\omega}$ is a unit vector describing the direction of the ray.
Light incident along $\mathbf{r}$ is integrated at a sensor pixel $(i,j)$ over time bins $n$, resulting in a transient measurement $\bm{\tilde{\tau}}[i,j,n]$. The number of photons collected in each time bin follows a Poisson distribution [29, 42]:
$\bm{\tilde{\tau}}[i,j,n]$ $\sim\text{Poisson}(N\eta\bm{\lambda}[i,j,n]+B),$
$\quad\text{where}\quad B$ $=N(\eta A[i,j]+D).$ (1)
Here, $\bm{\lambda}$ is the expected number of incident photons and depends on the reflectivity of the point imaged by pixel $(i,j)$. The value $N$ denotes the number of emitted laser pulses per pixel, $\eta$ is the detection efficiency of the sensor, and $B$ is the number of background detections, which dependens on the average ambient photon rate at pixel coordinates $(i,j)$, $A[i,j]$. Finally, the value $D$ is the dark count—i.e., the number of false detections produced by the sensor.
For a Poisson distribution, the signal-to-noise ratio (SNR) is defined as the ratio between the mean and its standard deviation. Hence, in our case, the SNR of $\tilde{\bm{\tau}}$ is given by:
$\frac{N\eta\bm{\lambda}[i,j,n]+B}{\sqrt{N\eta\bm{\lambda}[i,j,n]+B}}.$ (2)
Decreasing the number of laser pulses reduces the acquisition time, the number of collected photons, and the SNR of the measurements.
3.2 Surface-based Transient Rendering
We render transients by combining a neural representation that parameterizes the scene appearance and surface geometry with a time-resolved volume rendering equation.
Neural surface representation.
We use a neural surface representation $\mathcal{F}$ to represent the scene appearance and surface geometry. The representation consists of a multi-resolution hash-grid, a base multi-layer perceptron (MLP) $\mathcal{F}_{b}$, and a colour MLP $\mathcal{F}_{c}$ [35, 24]. The MLP $\mathcal{F}_{b}$ maps a point along a ray $\mathbf{r}(t)\in\mathbb{R}^{3}$ to a geometric latent feature vector $\bm{l}_{\text{geo}}\in\mathbb{R}^{16}$ and an SDF value $f\in\mathbb{R}$. The surface normal is given as $\bm{n}=\frac{\nabla f(\mathbf{r})}{\|\nabla f(\mathbf{r})\|}$, where the gradient $\nabla f(\mathbf{r})$ is computed via finite differences. We then recover the radiance at $\mathbf{r}(t)$ as $\bm{c}=\mathcal{F}_{\bm{c}}(\bm{n},\gamma(\bm{\omega}),\bm{l}_{\text{geo}})$, where $\gamma$ is a positional encoding function [47].
To make the representation compatible with time-resolved volume rendering, we follow NeuS [51] and convert the predicted SDF to a density field $\sigma(\mathbf{r})$ as
$\sigma(\mathbf{r})=\max\left(\frac{-\frac{d}{dt}\Phi_{s}\left(f(\mathbf{r})\right)}{\Phi_{s}\left(f(\mathbf{r})\right)},0\right),$ (3)
where $\Phi_{s}(x)=(1+e^{-sx})^{-1}$ is the sigmoid function, and $s$ is a learnable parameter controlling the support of the density field around the zero level set [53].
Time-resolved volume rendering.
Following TransientNeRF [29] we use the predicted density and radiance values from this representation to render transients $\bm{\tau}$ at each pixel $(i,j)$ and time bin $n$ using the time-resolved volume rendering equation:
$\bm{\tau}[i,j,n]$ $=\int_{\mathcal{P}_{i,j}}\int_{\mathcal{T}_{n}}(tc)^{-2}\,T(t)^{2}\sigma(\mathbf{r})\mathbf{c}(\mathbf{r},\bm{\omega})\,\mathrm{d}t\,\mathrm{d}\mathbf{p},$ (4)
$\text{where}\,\,\,T(t)$ $=\exp\left(-\int_{t_{0}}^{t}\sigma(\mathbf{r})\,\mathrm{d}s\right).$
The value $T(t)$ is the transmittance from a distance $t_{0}$ along the ray to $t$, and this term is squared to reflect the two-way propagation of light [1]. The value $(tc)^{-2}$ accounts for the inverse-square falloff of intensity, $\mathcal{T}_{n}$ is the transient bin width, and $\mathcal{P}_{i,j}$ models the pixel footprint. Note that in practice, we approximate the integral using quadrature by sampling at locations $t_{i}$ along the ray [30]. We approximate the integral over $\mathcal{P}_{i,j}$ by explicitly sampling multiple rays within the area of each pixel. Finally, we convolve $\bm{\tau}$ with the calibrated temporal impulse response of the lidar system to yield the final rendered transient $\bm{\tau}_{f}$ [29].
3.3 Optimization
Objective function.
As our main training objective, we supervise the rendered transient using the measurements $\mathbf{\tilde{\bm{\tau}}}$:
$\mathcal{L}_{\bm{\tau}}=\frac{1}{|\mathcal{R}_{\text{train}}|}\sum_{\mathcal{R_{\text{train}}}}\|\tilde{\bm{\tau}}-{\bm{\tau}_{f}}\|_{1},$ (5)
where $\mathcal{R}_{\text{train}}$ is the set of rays corresponding to the captured transient measurements used for training.
Weight variance regularization.
To recover thinner and smoother surfaces and aid the model in generalizing to viewpoints far from the training distribution, we regularize the scene density from sampled unseen viewpoints by penalizing the variance of the rendering weights along each ray.
To sample unseen viewpoints we follow RegNeRF [36] and sample camera origins on a sphere around the scene estimated from the training views (see supplement for details). We then regularize the density from these unseen views using RawNeRF’s weight-variance regularizer [31]. However, we find that replacing their rendered depth estimate with the argmax depth [29] helps recover more accurate surfaces, as shown in the supplement.
More specifically let $t_{i}$ be the samples along the ray used in the quadrature approximation of Eq. 4. Then the values $w_{i}=\sigma(t_{i})T(t_{i})$ represent alpha compositing weights applied to the radiance values along a ray. We use $w_{i}$ to compute the weight-variance regularizer, which we apply to rays $\mathcal{R}_{\text{train}}^{\prime}$ sampled from the set of unseen views:
$\mathcal{L}_{\text{weight\_var}}=\frac{1}{|\mathcal{R}_{\text{train}}^{\prime}|}\sum_{\mathcal{R}_{\text{train}}^{\prime}}\sum_{i}w_{i}\frac{(t_{i}-d)^{3}+(t_{i-1}-d)^{3}}{3(t_{i}-t_{i-1})}.$ (6)
Here, $d=\operatorname*{arg\,max}_{t^{*}}\;T(t^{*})\sigma(t^{*})$ represents the depth, or distance along the ray with the maximum probability of ray termination [29].
Reflectivity loss.
To improve performance in the low photon count regime, we introduce a reflectivity loss, which provides additional supervision using the integrated transients. Since the integrated transients have a higher signal-to-noise ratio than the individual photon count histogram bins, this provides a complementary source of supervision.
We implement this loss as
$\mathcal{L}_{\text{ref}}=\frac{1}{|\mathcal{R}_{\text{train}}|}\sum_{\mathcal{R_{\text{train}}}}\|\sum_{n}\bm{\tau}[i,j,n]-\sum_{n}\tilde{\bm{\tau}}[i,j,n]\|_{1}.$ (7)
In summary, our training loss can be expressed as the following:
$\mathcal{L}$ $=\mathcal{L}_{\bm{\tau}}+\lambda_{\text{ref}}\mathcal{L}_{\text{ref}}+\lambda_{\text{sc}}\mathcal{L}_{\text{sc}}+\lambda_{\text{eik}}\mathcal{L}_{\text{eik}}$
$\quad+\lambda_{\text{weight\_var}}\mathcal{L}_{\text{weight\_var}}+\lambda_{\text{sparse}}\mathcal{L}_{\text{sparse}},$ (8)
where
$\mathcal{L}_{\text{sc}}=\sum\limits_{\begin{subarray}{c}i,j,n\\ \bm{\widetilde{\tau}}[i,j,n]<B\end{subarray}}\int_{\mathcal{P}_{i,j}}\int_{\mathcal{T}_{n}}T(t)\sigma(\mathbf{r})\,\mathrm{d}t\,\mathrm{d}\mathbf{p},$ (9)
is the space carving loss [29], which penalizes densities at locations where the transient is below a background level $B$.
$\mathcal{L}_{eik}=\frac{1}{|\mathbf{r}\in\mathcal{R}_{\text{train}}|}\sum_{\mathbf{r}\in\mathcal{R}_{\text{train}}}(\|\nabla f(\mathbf{r})\|-1)^{2},$ (10)
is the Eikonal regularization [12], which encourages the SDF to have gradient magnitude equal to one, and
$\mathcal{L}_{\text{sparse}}=\frac{1}{|\mathbf{r}\in\mathcal{R}_{\text{train}}|}\sum_{\mathbf{r}\in\mathcal{R}_{\text{train}}}\exp(-\alpha|f(\mathbf{r})|),$ (11)
is the sparsity loss, which helps to prevent floating artifacts, and $\alpha$ is a scalar hyperparameter.
3.4 Implementation Details
We implement our method in Pytorch-Lightning [8, 13]. We use the NerfAcc [23] version of InstantNGP [35] for efficient ray marching, and we build on top of TransientNeRF’s [29] time-resolved rendering.
Training and optimization.
We initialize the SDF to approximate a sphere [2]. We follow Neuralangelo’s coarse-to-fine optimization strategy by initializing the step-size in the gradient calculation $\epsilon$ as the coarsest hash-grid voxel size and exponentially decreasing it during training. For the hash-grid, we start at 4 hash resolutions and unmask two new hash resolutions every 5000 steps. We optimize the network using the AdamW [27] optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and employ a two-step learning rate scheduler by initializing the learning rate to $1\times 10^{-5}$ and linearly increasing it until it reaches $1\times 10^{-3}$ by the end of the 5000 steps. Then, the learning rate is exponentially decreased over the remaining steps, with a total of 250K steps. All experiments were run on a single Nvidia RTX A6000 GPU with 48GB of memory. We provide additional implementation details, including hyperparameter settings, in the supplement.
4 Multiview Transient Dataset
To test the robustness of different methods on low-SNR transient data we adapt the simulated and captured datasets introduced by TransientNeRF [29]. For each captured and simulated scene, we synthesize transients with an average of 300, 150, 50 and 10 photons per occupied (non-background) pixel. The paragraphs below outline how these datasets are created.
Simulated dataset.
To synthesize transients at different photon counts in simulation we match the flux from the time-resolved Mitsuba renderer [29] to the target number of photons per pixel. We then follow Malik et al. [29] and set the background flux to $0.1$ photons per $2850$ captured scene photons, which is set to match the statistics of the captured dataset. The flux is then Poisson sampled to simulate photon counts.
Captured dataset.
We create low-flux captured datasets using the raw photon arrival timestamps from TransientNeRF. Specifically, we subsample the single photon arrivals using thinning [22] to achieve a set number of detected photons. We then bin the photon arrivals into a transient histogram. Additional dataset details can be found in the supplement. Both the simulated and captured variable photon count datasets will be released upon publication.
5 Results
Figure 3: Results on the simulated dataset using an average of 1500 photons per occupied (non-background) pixel. We show the recovered meshes from the baselines and the proposed method. For the transient-based methods, we also show rendered transients for the indicated pixels. Our method recovers smoother meshes with fewer missing parts. We also recover transients that better match the ground truth.
Table 1: Simulated results assessing image quality, depth accuracy and mesh quality. The left column indicates average photons per pixel.
Chamfer Distance \,$\downarrow$ PSNR (dB)\,$\uparrow$ LPIPS\,$\downarrow$ L1 (depth) \,$\downarrow$ Transient IOU\,$\uparrow$
Method 2 views 3 views 5 views 2 views 3 views 5 views 2 views 3 views 5 views 2 views 3 views 5 views 2 views 3 views 5 views
6000 Neuralangelo [24] 4.56 7.99 7.17 18.57 19.44 22.03 0.388 0.442 0.434 0.162 0.122 0.090 - - -
RegNeRF  [36] 4.94 3.48 4.38 18.32 19.89 21.67 0.376 0.412 0.418 0.264 0.142 0.100 - - -
MonoSDF w/ mask  [63] 3.78 1.67 0.18 19.76 20.74 24.65 0.240 0.244 0.132 0.045 0.026 0.015 - - -
TransientNeRF [29] 0.27 0.44 0.31 21.38 23.48 28.39 0.172 0.151 0.115 0.015 0.011 0.013 0.31 0.40 0.55
Proposed 0.07 0.06 0.05 25.28 26.69 29.49 0.176 0.155 0.150 0.006 0.006 0.005 0.54 0.58 0.68
300 TransientNeRF 0.26 0.33 0.30 21.34 22.79 25.88 0.201 0.152 0.121 0.012 0.012 0.015 0.32 0.41 0.58
Proposed 0.07 0.06 0.06 25.70 26.65 28.50 0.196 0.162 0.149 0.007 0.007 0.005 0.56 0.60 0.70
150 TransientNeRF 0.25 0.65 0.27 21.04 22.18 25.35 0.203 0.161 0.117 0.015 0.013 0.010 0.31 0.39 0.56
Proposed 0.09 0.07 0.29 25.56 26.52 28.19 0.193 0.165 0.151 0.007 0.006 0.005 0.56 0.60 0.70
50 TransientNeRF 0.53 0.66 0.50 20.16 21.57 24.01 0.232 0.184 0.133 0.021 0.017 0.014 0.29 0.40 0.55
Proposed 0.15 0.08 0.06 24.73 25.76 26.55 0.218 0.163 0.167 0.012 0.005 0.007 0.53 0.59 0.63
10 TransientNeRF 2.86 0.31 5.50 18.94 22.85 16.73 0.265 0.218 0.209 0.052 0.018 0.092 0.27 0.43 0.19
Proposed 0.22 0.28 0.18 22.76 23.98 25.51 0.254 0.216 0.208 0.016 0.015 0.014 0.43 0.47 0.54
Figure 4: Results on the captured dataset. We show the recovered meshes from the baselines and the proposed method. Due to the lack of ground-truth, we include the closest captured image for reference of the scene. As can be seen, Neuralangelo recovers smoother meshes with fewer missing parts. We also recover transients that better match the ground truth.
Table 2: Captured results assessing image quality, depth accuracy and mesh quality. The left column indicates average photons per pixel.
PSNR (dB)\,$\uparrow$ LPIPS\,$\downarrow$ L1 (depth) \,$\downarrow$ Transient IOU\,$\uparrow$
Method 2 views 3 views 5 views 2 views 3 views 5 views 2 views 3 views 5 views 2 views 3 views 5 views
1500 Neuralangelo [24] 17.79 19.28 21.07 0.380 0.355 0.300 0.076 0.071 0.061 - - -
RegNeRF  [36] 18.50 19.59 20.29 0.375 0.320 0.330 0.085 0.073 0.076 - - -
MonoSDF w/ mask [45] 17.63 21.16 27.25 0.353 0.248 0.157 0.033 0.026 0.021 - - -
TransientNeRF [29] 22.11 21.83 22.72 0.271 0.212 0.172 0.005 0.006 0.010 0.34 0.41 0.50
Proposed 21.31 23.95 25.22 0.241 0.170 0.145 0.007 0.006 0.006 0.40 0.50 0.58
300 TransientNeRF 20.14 21.99 20.53 0.270 0.210 0.177 0.012 0.007 0.014 0.26 0.39 0.36
Proposed 21.60 23.53 23.79 0.223 0.170 0.151 0.005 0.005 0.012 0.42 0.49 0.55
150 TransientNeRF 19.13 21.06 19.95 0.277 0.206 0.196 0.016 0.015 0.018 0.21 0.33 0.28
Proposed 21.23 23.56 25.33 0.222 0.171 0.146 0.007 0.005 0.004 0.41 0.50 0.60
50 TransientNeRF 18.80 19.63 19.31 0.274 0.194 0.190 0.020 0.020 0.018 0.19 0.26 0.24
Proposed 19.14 22.33 25.34 0.280 0.180 0.148 0.017 0.011 0.007 0.34 0.44 0.59
10 TransientNeRF 16.40 16.36 15.40 0.246 0.259 0.272 0.037 0.041 0.090 0.08 0.08 0.02
Proposed 22.40 23.05 25.47 0.242 0.208 0.173 0.008 0.008 0.007 0.35 0.36 0.45
We evaluate our method for transient and intensity image synthesis from novel viewpoints, depth rendering, and mesh recovery. We also benchmark the performance of our method’s robustness to low-SNR data. A more comprehensive set of results and ablation studies can be found in the supplement.
Baselines.
We compare the rendered intensity, depth maps, and meshes against state-of-the-art methods in surface reconstruction and few-view surface reconstruction: Neuralangelo [24], RegNeRF [36], and MonoSDF [63]. To ensure a fair comparison, we implement these baselines using the same Neuralangelo [24] backbone architecture, and we integrate their proposed loss functions. All three methods are supervised using integrated transient measurements. MonoSDF is also supervised using depth maps obtained by applying a log-matched filter [18] to the transient measurements, and the input normal maps are estimated from the depth. We also compare our method to TransientNeRF [29].
Evaluation criteria.
To measure the mesh reconstruction quality we use Chamfer Distance. For image reconstruction evaluation we use PSNR and LPIPS [64]. We also evaluate the quality of the rendered transients using Transient IOU [29], defined as the intersection over the union of the rendered transient and the ground truth transient (see supplement for details).
5.1 Simulated Results
Figure 3 shows the recovered meshes and rendered transients for specific pixels from the proposed method compared to baselines. We show results for measurements with an average of 1500 photons per occupied pixel (results for other photon levels are included in the supplement). Due to a lack of any depth-based priors, Neuralangelo [24] and RegNeRF [36] both fail to recover accurate scene geometry. MonoSDF [63] recovers noisier surfaces, than the proposed approach, especially when trained with fewer input viewpoints. Compared to TransientNeRF [29], the proposed method recovers smoother surfaces with fewer missing parts and fewer floating artifacts, especially in the case of training with two or three input viewpoints. The results highlight the benefit of training a neural surface representation with transient supervision.
Figure 5: Novel view synthesis for varying photon levels. We show the rendered novel view on the cinema scene, trained on five viewpoints with an average of $50,300$ photons. We also show peak-time visualizations [28], which show the full transient in a single visualization. Hue encodes the time of peak intensity, brightness is modulated by the maximum intensity, and each band corresponds to an isochrone, or wavefront of equal path length. We show transient plots (right) for the pixel indicated on the ground-truth image.
Table 1 shows the quantitative performance of the proposed method against the baselines averaged across all scenes in the two, three, and five training viewpoint settings. We benchmark performance for varying measurement SNRs by training using transients with lower photon counts. To compute the metrics, we first scale the rendered transients so that the average photon count per occupied pixel (across all viewpoints) is the same across all photon levels.
The recovered meshes from the proposed method achieve significantly lower Chamfer Distance compared to the other methods and roughly 5×5$\times$5 × lower than the second-best method, TransientNeRF. The recovered depth maps are also more accurate than the baseline methods, and the image metrics are comparable. Finally, the transients rendered from our method are more similar to the ground truth based on the Transient IOU metric.
Table 3: Ablation study on the proposed method in the Lego scene. We present results without the sparsity loss (w/o $\mathcal{L}_{\text{sparse}}$), without the space carving loss (w/o $\mathcal{L}_{\text{sc}}$) and without the weight variance regularizer (w/o $\mathcal{L}_{\text{weight\_var}}$).
Chamfer Distance \,$\downarrow$ PSNR (dB)\,$\uparrow$ Transient IOU\,$\uparrow$
Method 2 views 3 views 5 views 2 views 3 views 5 views 2 views 3 views 5 views
Ours w/o $\mathcal{L}_{\text{sc}}$ 0.07 0.04 0.11 25.63 26.32 26.22 0.53 0.57 0.53
Ours w/o $\mathcal{L}_{\text{weight\_var}}$ 0.28 0.37 1.04 21.58 22.66 23.81 0.48 0.53 0.60
Ours w/o $\mathcal{L}_{\text{sparse}}$ 0.02 0.03 0.08 25.43 26.87 26.58 0.53 0.58 0.63
Ours 0.02 0.02 0.01 25.72 26.66 27.47 0.53 0.58 0.63
5.2 Captured Results
Figure 4 shows the recovered meshes for our method and the baselines on the food, carving, cinema scenes for two, three, and five training viewpoints respectively (see the supplement for qualitative results for all scenes). As in TransientNeRF [29], we mask out the background of the transients for the captured dataset. Similarly to the simulated dataset, the meshes recovered from the proposed method are both smoother and contain fewer artifacts than the baseline methods.
We report metrics averaged across six captured scenes in Table 2. Due to a lack of ground-truth meshes, we are not able to report Chamfer distances and the depth L1 metric is calculated by comparing to depth maps estimated from log-matched filtering of transients from held out views. Our quantitative results show an improvement over TransientNeRF across all image and depth metrics. Overall, our method remains robust across various photon levels, whereas TransientNeRF’s image rendering quality deteriorates rapidly as the mean photon count decreases.
Figure 5 shows the results of training our method and TransientNeRF [29] on transients with an average of $50$ and $300$ photons per occupied pixel. At these photon levels, TransientNeRF rendered transients’ intensities are weaker than the measured ones. On the other hand, due to the reflectivity loss (see supplement), our method recovers more accurate transients.
5.3 Ablation Study
Table 3 ablates the effect of the regularization terms on the simulated Lego scene. The results show that all the added regularizers are vital to the performance; however, the weight-variance regularizer has the greatest effect. Applying this regularizer to the rendered unseen viewpoints is especially important in the few-viewpoint setting.
6 Discussion
We show that using a neural surface representation with transient supervision and additional regularization recovers high-fidelity surfaces in the few viewpoint setting with as few as 10 measured photons per pixel. Our work improves the practicality of transient-based reconstruction by reducing the required light levels, which could be especially useful for 3D imaging of photo-sensitive materials or long-range (and low-SNR) remote sensing.
There are multiple limitations of our method that could be addressed in future work. For example, computational efficiency could potentially be improved by developing new time-resolved rendering frameworks based on Gaussian splatting [16]. While the main focus of our work is scene reconstruction using the direct component of light, transients carry much richer scene information, including higher order bounces of light. Future work could investigate such techniques for systems that infer further scene properties such as reflectance, albedo, or other material properties.